OpenStudy (anonymous):

Use the fundamental Theorem of Algebra to determine the number of zeros. Then find the zeros. f(x)= (x^2 - 2x - 24)(x^2 + 4)

OpenStudy (anonymous):

The Fundamental Theorem of Algebra says that a polynomial has a number of roots equal to the degree of the polynomial. To determine the degree of the polynomial, let's multiply the two quantities to get \[x^4 - 2x^3 - 20x^2 - 8x - 96\]What is the degree of this polynomial? (The degree is the value of the highest exponent.)

OpenStudy (anonymous):

The degree is 4, which means there are 4 zeros? and then how do you find the zeros from there?

OpenStudy (anonymous):

Correct! We need to set the function equal to zero and solve for x \[x^4 - 2x^3 - 20x^2 - 8x - 96 = 0\] To make things easier, lets return to our original state. Now we have two 2nd order equations to solve. We get \[x^2 - 2x - 24 = 0 ~~\text{and}~~ x^2 + 4 = 0\]

OpenStudy (anonymous):

(x-6)(x+4)=0 and (x+2)(x-2)=0 so the zeros are 6, -4, 2, and -2?

OpenStudy (anonymous):

Correct. Note that only imaginary roots are present in the x^2 + 4 equation.

OpenStudy (anonymous):

Wait, why are there only imaginary roots present in x^2 + 4?

OpenStudy (anonymous):

x^2 = -4. You can't take the square root of a negative number.

OpenStudy (anonymous):

Can you not factor the equation so that it's (x+2)(x-2) and get the answer of +-2? If not, how come?

OpenStudy (anonymous):

You can. Your answers are correct. I was just noting the fact that they were imaginary.